Melissa's brother is 7 years old. Three years ago she was as old as he is now. How old will Melissa be in 4 years?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

----------------------------------------
Melissa's brother is 7 years old.
---------------------------------------
call b, the age for Melissa's brother.
b=7.


--------------------------------------------
Three years ago she was as old as he is now.
-------------------------------------------

m, for Melissa.
m-3=7
-
m=10----------Melissa is currently 10 years of age.



----------------------------------------------
How old will Melissa be in 4 years?
----------------------------------------------

Answer by ikleyn(52781)

.
Three years ago, she was 7 years old - hence, now she is 10 years old, and in 4 years she will be 14 years old.


Second grade level, when they even do not know what an equation is . . . and how to solve them . . .